# On elliptic and quasiregularly elliptic manifolds

### v1

@@ 4. Open surfaces
The main difficulty in analyzing such examples is determining whether they are Gromov-elliptic. To determine whether a surface is quasiregularly elliptic, we use a criterion due to Ahlfors [1]: a metric on the plane is parabolic if and only if $\int_{0}^{\infty}\frac{dr}{L(r)}$ diverges, where $L(r)$ is the circumference of the geodesic circle of radius $r$ centered at some fixed point. A special case of this was highlighted and given a simple proof by Milnor [25]: a rotationally symmetric metric is parabolic if its curvature $K$ satisfies $K\geq-\frac{1}{r^{2}\log r}$ for large radii $r$, and hyperbolic if it satisfies $K\leq-\frac{1+\varepsilon}{r^{2}\log r}$.
@@ 4.1. A Gromov-elliptic surface which is conformally hyperbolic
Fix a geodesic $\gamma$ starting at the origin, for example the $y$-axis. We first understand the exponential map $\varphi_{R}$ at a point $x=\gamma(R)$, applied to a disk of radius $R/2$ around the origin. The curvature on the image of this map is pinched between $-4R^{-2}$ and $-R^{-2}/4$. Then by Toponogov’s theorem, the exponential map from a Euclidean disk of radius $R/2$ to this disk is $4\sinh(1)$-Lipschitz. On the other hand, the map clearly expands areas everywhere.
@@ 4.2. A parabolic surface which is not Gromov-elliptic
where $\varphi:[0,1]\to[0,1]$ is a smooth, non-increasing bump function supported on $[0,\frac{1}{2})$ and equal to $1$ on $[0,\frac{1}{4}]$. Then $\Sigma$ is parabolic, but not Gromov-elliptic.
@@
Next we push the curve down the side of the spike, as with a squeegee: a point starts moving once the homotopy reaches down to its height. This homotopy takes time $<\log(2R)+\frac{1}{4}<\log R+1$ and is $\pi$-Lipschitz, since distances increase by a factor of at most $2$ from $g$.
@@
Consider the homology class induced by $f(\partial B)$ in $G^{(1)}$. This induces a unique cellular cycle $z\in C_{1}(G)$; the mass of $z$ (that is, the sum of the coefficients) is $O(\operatorname{length}(f(\partial B)))=O(R)$. Let $c$ be the (likewise unique) $2$-chain filling $z$; we must estimate the portion of the mass of $c$ contained in $[-\sqrt{R},\sqrt{R}]^{2}$.
@@
Now let $V$ be the portion of the mass of $c$ contained in $[-\sqrt{R},\sqrt{R}]^{2}$. We have
$V\leq\sum_{i=1}^{\infty}\min\{a(i),4R\}\leq 2\sqrt{R}\sum_{i=1}^{\infty}\sqrt{a(i)}\leq\frac{1}{2}\sqrt{R}\operatorname{mass}(z)=O(R^{3/2}).\qed$
@@
#(left turns)#(right $\frac{\#(\text{left turns})-\#(\text{right turns})}{4}.$
@@
Now we slightly modify the concept of turning number to be well-defined for free homotopy classes of curves in $G^{(1)}\cup[-\sqrt{R},\sqrt{R}]^{2}$, which we denote by $t_{\operatorname{sq}}$. Here, we produce a canonical representative by replacing each section of curve that traverses $[-\sqrt{R},\sqrt{R}]^{2}$ by any simple curve between its endpoints. It’s easy to see that this gives rise to a well-defined notion of turning number and that Lemma 4.5 still holds for $t_{\operatorname{sq}}$.
@@
by Lemma 4.5(a). Each nontrivial component winds around a region which contains a preimage $u\in f^{-1}(a+\frac{1}{2},b+\frac{1}{2})$. Since $f$ is $1$-Lipschitz, the region contains a ball of radius $\frac{1}{2}\log R$ around $u$. Therefore, the number of nontrivial components #$\#\{\gamma_{i}\}$ is $O(R^{2}/\log^{2}R)$.
Since $f$ is $1$-Lipschitz, $f_{\partial B}$ has length $O(R)$, so by Lemma 4.5(b), $t_{\operatorname{sq}}(f|_{\partial B})=O(R)$. It follows that $\sum_{i}t(\gamma_{i})=O(R^{2}/\log^{2}R)$. This gives the sum of local degrees of $f$ over points not in $[-\sqrt{R},\sqrt{R}]^{2}$. Combining this with Lemma 4.4, which covers the points inside this square, we get the result. ∎
Grid squares in the image of $f$ have area bounded by $14\log((\pi\sqrt{2}+1)R)+$ 1. Moreover, since we assumed that $f|_{\partial B}$ maps to $G^{(1)}$, $f$ has the same local degree over all points in a grid square. By Lemma 4.6,
@@ 4.3. Some remarks on surfaces with nontrivial topology
which maps the boundary of the domain to the boundary of the disk. This map is defined explicitly on $[0,1]\times[0,\infty)$ by the polar coordinate formula
@@ 5. General open manifolds
Now we expand the discussion from surfaces to open manifolds of all dimension. Here we give further constructions of manifolds which are elliptic, but not quasiregularly elliptic. For example, we show that every open manifold has an elliptic metric, and at least a large class (all manifolds with a spherical end) have elliptic metrics of finite volume. A great many of these manifolds admit no quasiregularly elliptic metric. We also prove Theorem 1.9, which proves a restriction on the fundamental group of an open elliptic manifold whose metric satisfies a certain rather strong assumption; this is analogous to a result for quasiregular ellipticity which requires no such assumption. Nevertheless, even open subsets of compact manifolds (for which Theorem 1.9 holds) may be elliptic without being quasiregularly elliptic:
@@
Let $n\geq 2$, and consider $S^{n-2}$ embedded in the round $S^{n}$ as the intersection with an $(n-1)$-plane through the origin in $\mathbb{R}^{n+1}$. The open manifold $S^{n}\setminus S^{n-2}$ is Gromov-elliptic, and therefore so is $S^{n}\setminus X$ for any $X\subseteq S^{n-2}$.
In contrast, for any countably infinite set $X$, a generalization of the Rickman–Picard theorem [34] by Holopainen and Rickman [18, Theorem 3.1] shows that $S^{n}\setminus X$ does not admit a metric under which it is quasiregularly elliptic.
@@ 5.1. An elliptic metric on any open manifold
Let $\gamma:[0,\infty)\to M$ be a ray, that is, a proper embedding parametrized by arclength. We modify $(M,g)$ by cutting out a small tubular neighborhood of $\gamma([0,\infty))$ and replacing it with an isometric copy of $\mathbb{R}^{n}$ missing a small tubular neighborhood of the negative $x_{1}$-axis. After a bit of smoothing, we get a new Riemannian manifold $(M^{\prime},g^{\prime})$ which is obviously diffeomorphic to $M$ and which includes an isometric copy of $[1,\infty)\times\mathbb{R}^{n-1}$. But then the map
@@ 5.2. A finite-volume elliptic manifold with any fundamental group
Specifically, we build a metric $\tilde{g}$ on $C=S^{n-1}\times[0,\infty)$ by embedding it in $\mathbb{R}^{n+1}$ as a solid of revolution obtained by revolving the graph of a function $\rho(t)$. We define $\rho$ so that for every $p\in\mathbb{N}$, it has a local minimum $\rho(p)=2^{-\frac{2p}{n-1}}$ at $p$, and a local maximum $\rho(p+1/2)=2^{-\frac{p}{n-1}}$; between these, the value of $\rho$ is interpolated smoothly so that the average value over the interval is proportional to the local maximum. In particular, each “nodule” $N_{p}$ from $p$ to $p+1$ has volume proportional to $2^{-p}$, and the total volume of $(C,\tilde{g})$ is finite. In between the nodules, at $p$, the metric has a spherical neck of surface area proportional to $2^{-2p}$.
We now describe a wrapping map $f:\mathbb{R}^{n}\to(C,\tilde{g})$. It will be defined on a sequence of cubes
@@
Inside $Q_{p}$, we choose a submanifold $S$ isometric to a $(n-1)$-dimensional rectangle which is folded up like an accordion, with $\partial S\subseteq\partial Q_{p}$, and so that the exponential map on its normal $1$-disk bundle is an embedding and fills up a proportion of $Q_{p}$ which depends only on $n$. Restricting this exponential map $S\times[-\frac{1}{2},\frac{1}{2}]$ gives a codimension 0 submanifold $Q_{p}^{0}\subseteq Q_{p}$ which is $2$-bilipschitz to $S\times[-\frac{1}{2},\frac{1}{2}]$ and separates the rest of $Q_{p}$ into two components $Q_{p}^{+}$ (above) and $Q_{p}^{-}$ (below).
@@
Note that $\pi_{1}(M)$ may be an infinitely generated group, for example $\mathbb{Q}$. The hypothesis holds for example for manifolds of pinched negative curvature as well as real analytic manifolds of nonpositive curvature, including locally symmetric spaces. The proof in both cases is the same and follows from results described in [4, §10, 13].
@@
$h(M)=\inf\left\{\frac{\operatorname{vol}(\partial A)}{\operatorname{vol}A}:A\text{ is a relatively compact open subset with }\operatorname{vol}A\leq\frac{1}{2}\operatorname{vol}M\right\}.$
@@
The Cheeger constant of a manifold can be zero, for example for those constructed in Section 5.2. In future work, we will show that $M$ has nonzero Cheeger constant if and only if it satisfies the hypothesis in Theorem 1.9. This equivalence has an analogue for manifolds of infinite volume stated by Gromov [12, §4.1] which can be proved using similar techniques.
@@
Let $f:\mathbb{R}^{n}\to M$ be a wrapping map.
By Theorem LABEL:IP-dual, we can write $d\operatorname{vol}_{M}=\omega+d\alpha$, where $\omega$ is compactly supported and $\alpha$ is bounded. Then by Stokes’ theorem,
@@
Now let $K^{\prime}\supset K$ be another compact set which contains the $r$-neighborhood of $K$; we can assume $K^{\prime}$ is a smooth codimension zero submanifold of $M$ and give it a triangulation $\mathcal{T}$. Let $\lambda>0$ be such that every $\lambda$-ball in $K^{\prime}$ is contained in the star of some vertex of $\mathcal{T}$. Choose a triangulation $\mathcal{T}^{\prime}$ of $B(R)$ whose simplices are $C(n)$-bilipschitz to a linear simplex of side length $\min\{r,\lambda\}$ and have diameter $<\min\{r/2,\lambda/2\}$. Then there is a subcomplex $X\subseteq\mathcal{T}^{\prime}$ which contains the image of $K$ and whose star lies in $K^{\prime}$. Moreover, by the usual proof of the simplicial approximation theorem, see e.g. [14], a perturbation of $f$ gives a map $f^{\prime}$ whose restriction to $X$ is simplicial on $\mathcal{T}$.
We can assume that the induced map $\pi_{1}(K)\to\pi_{1}(M)$ is injective. Now consider the induced map from the dual graph of $X$ to the dual graph $\Gamma$ of $\widetilde{T}$, the universal cover of $T$. This defines a 00-cochain in $\Gamma$ of volume $\sim R^{n}$ whose coboundary has volume $\sim R^{n-1}$. Since $\Gamma$ is quasi-isometric to the Cayley graph of $\pi_{1}(K)$, it has the same isoperimetric and growth functions. Therefore, by the Varopoulos isoperimetric inequality (see e.g. [13, §6.E]), $\pi_{1}(K)$ has growth rate $O(R^{n})$.
Now we could have taken $\pi_{1}(K)$ to be any finitely generated subgroup of $\pi_{1}(M)$. Therefore the nilpotency class of any such subgroup is at most $n-1$ (in fact, at most around $\sqrt{n}$, but this is harder to see), and so is the nilpotency class of $\pi_{1}(M)$. ∎
@@
First define a Lipschitz map of positive asymptotic degree $g_{n-2}:\mathbb{R}^{n-2}\to S^{n-2}$; for example, we can compose the universal covering map $\mathbb{R}^{n-2}\to T^{n-2}$ with a map that collapses the $(n-3)$-skeleton of the torus.
@@ Acknowledgements
The authors would like to thank Ben Lowe for comments drawing attention to the parallels between their previous work, which sparked this collaboration. F.M. would also like to thank Pekka Pankka and Susanna Heikkilä for useful conversations and comments.

### v3

@@ 5. Open surfaces
The main difficulty in analyzing such examples is determining whether they are Gromov-elliptic. To determine whether a surface is quasiregularly elliptic, we use a criterion due to Ahlfors [Ahl]: a metric on the plane is parabolic if and only if $\int_{0}^{\infty}\frac{dr}{L(r)}$ diverges, where $L(r)$ is the circumference of the geodesic circle of radius $r$ centered at some fixed point. A special case of this was highlighted and given a simple proof by Milnor [Mil]: a rotationally symmetric metric is parabolic if its curvature $K$ satisfies $K\geq-\frac{1}{r^{2}\log r}$ for large radii $r$, and hyperbolic if it satisfies $K\leq-\frac{1+\varepsilon}{r^{2}\log r}$.
@@ 5.1. A Gromov-elliptic surface which is conformally hyperbolic
Fix a geodesic $\gamma$ starting at the origin, for example the $y$-axis. We first understand the exponential map $\varphi_{R}$ at a point $x=\gamma(R)$, applied to a disk of radius $R/2$ around the origin. The curvature on the image of this map is pinched between $4R^{-2}$ and $R^{-2}/4$. Then by Toponogov’s theorem, the exponential map from a Euclidean disk of radius $R/2$ to this disk is $4\sinh(1)$-Lipschitz. On the other hand, the map clearly expands areas everywhere.
@@ 5.2. A parabolic surface which is not Gromov-elliptic
where $\varphi:[0,1]\to[0,1]$ is a smooth, non-increasing bump function supported on $0,\frac{1}{2})$ and equal to $1$ on $0,\frac{1}{4}]$. Then $\Sigma$ is parabolic, but not Gromov-elliptic.
@@
Next we push the curve down the side of the spike, as with a squeegee: a point starts moving once the homotopy reaches down to its height. This homotopy takes time $\log(2R)+\frac{1}{4}<\log R+1$ and is $\pi$-Lipschitz, since distances increase by a factor of at most $2$ from $g$.
@@
Consider the homology class induced by $f(\partial B)$ in $G^{(1)}$. This induces a unique cellular cycle $z\in C_{1}(G)$; the mass of $z$ (that is, the sum of the coefficients) is $O(\operatorname{length}(f(\partial B)))=O(R)$. Let $c$ be the (likewise unique) $2$-chain filling $z$; we must estimate the portion of the mass of $c$ contained in $\sqrt{R},\sqrt{R}]^{2}$.
@@
Now let $V$ be the portion of the mass of $c$ contained in $\sqrt{R},\sqrt{R}]^{2}$. We have
$V\leq\sum_{i=1}^{\infty}\min\{c(i),4R\}\leq 2\sqrt{R}\sum_{i=1}^{\infty}\sqrt{c(i)}\leq\frac{1}{2}\sqrt{R}\operatorname{mass}(z)=O(R^{3/2}).\qed$
@@
$\frac{\#(\text{left turns})-\#(\text{right turns})}{4}.$
@@
Now we slightly modify the concept of turning number to be well-defined for free homotopy classes of curves in $G^{(1)}\cup[-\sqrt{R},\sqrt{R}]^{2}$, which we denote by $t_{\operatorname{sq}}$. Here, we produce a canonical representative by replacing each section of curve that traverses $\sqrt{R},\sqrt{R}]^{2}$ by any simple curve between its endpoints. It’s easy to see that this gives rise to a well-defined notion of turning number and that Lemma 5.5 still holds for $t_{\operatorname{sq}}$.
@@
by Lemma 5.5(a). Each nontrivial component winds around a region which contains a preimage $u\in f^{-1}(a+\frac{1}{2},b+\frac{1}{2})$. Since $f$ is $1$-Lipschitz, the region contains a ball of radius $\frac{1}{2}\log R$ around $u$. Therefore, the number of nontrivial components $\{\gamma_{i}\}$ is $O(R^{2}/\log^{2}R)$.
Since $f$ is $1$-Lipschitz, $f_{\partial B}$ has length $O(R)$, so by Lemma 5.5(b), $t_{\operatorname{sq}}(f|_{\partial B})=O(R)$. It follows that $\sum_{i}t(\gamma_{i})=O(R^{2}/\log^{2}R)$. This gives the sum of local degrees of $f$ over points not in $\sqrt{R},\sqrt{R}]^{2}$. Combining this with Lemma 5.4, which covers the points inside this square, we get the result. ∎
Grid squares in the image of $f$ have area bounded by $4\log((\pi\sqrt{2}+1)R)+1$. Moreover, since we assumed that $f|_{\partial B}$ maps to $G^{(1)}$, $f$ has the same local degree over all points in a grid square. By Lemma 5.6,
@@ 5.3. Some remarks on surfaces with nontrivial topology
which maps the boundary of the domain to the boundary of the disk. This map is defined explicitly on $0,1]\times[0,\infty)$ by the polar coordinate formula
@@ 6. General open manifolds
Now we expand the discussion from surfaces to open manifolds of all dimension. Here we give further constructions of manifolds which are elliptic, but not quasiregularly elliptic. For example, we show that every open manifold has an elliptic metric, and at least a large class (all manifolds with a spherical end) have elliptic metrics of finite volume. A great many of these manifolds admit no quasiregularly elliptic metric. We also prove Theorem 1.10, which gives a restriction on the fundamental group of an open elliptic manifold whose metric satisfies a certain rather strong assumption; this is analogous to a result for quasiregular ellipticity which requires no such assumption. Nevertheless, even open subsets of compact manifolds (for which Theorem 1.10 holds) may be elliptic without being quasiregularly elliptic:
@@
Let $n\geq 2$, and consider $S^{n-2}$ embedded in the round $S^{n}$ as the intersection with an (n1)(n-1)-plane through the origin in $\mathbb{R}^{n+1}$. The open manifold $S^{n}\setminus S^{n-2}$ is Gromov-elliptic, and therefore so is $S^{n}\setminus X$ for any $X\subseteq S^{n-2}$.
In contrast, for any countably infinite set $X$, a generalization of the Rickman–Picard theorem [Rick] by Holopainen and Rickman [HoR2, Theorem 3.1] shows that $S^{n}\setminus X$ does not admit a metric under which it is quasiregularly elliptic.
@@ 6.1. An elliptic metric on any open manifold
Let $\gamma:[0,\infty)\to M$ be a ray, that is, a proper embedding parametrized by arclength. We modify $(M,g)$ by cutting out a small tubular neighborhood of $\gamma([0,\infty))$ and replacing it with an isometric copy of $\mathbb{R}^{n}$ missing a small tubular neighborhood of the negative $x_{1}$-axis. After a bit of smoothing, we get a new Riemannian manifold $(M^{\prime},g^{\prime})$ which is obviously diffeomorphic to $M$ and which includes an isometric copy of $1,\infty)\times\mathbb{R}^{n-1}$. But then the map
@@ 6.2. A finite-volume elliptic manifold with any fundamental group
[…]
@@
Inside $Q_{p}$, we choose a submanifold $S$ isometric to a (n1)(n-1)-dimensional rectangle which is folded up like an accordion, with $\partial S\subseteq\partial Q_{p}$, and so that the exponential map on its normal $1$-disk bundle is an embedding and fills up a proportion of $Q_{p}$ which depends only on $n$. Restricting this exponential map $S\times[-\frac{1}{2},\frac{1}{2}]$ gives a codimension 0 submanifold $Q_{p}^{0}\subseteq Q_{p}$ which is $2$-bilipschitz to $S\times[-\frac{1}{2},\frac{1}{2}]$ and separates the rest of $Q_{p}$ into two components $Q_{p}^{+}$ (above) and $Q_{p}^{-}$ (below).
@@
Note that $\pi_{1}(M)$ may be an infinitely generated group, for example $\mathbb{Q}$. The hypothesis holds for example for manifolds of pinched negative curvature as well as real analytic manifolds of nonpositive curvature, including locally symmetric spaces. The proof in both cases is the same and follows from results described in [BGS, §10, 13].
@@
[…]
@@
The Cheeger constant of a manifold can be zero, for example for those constructed in Section 6.2. In future work, we will show that $M$ has nonzero Cheeger constant if and only if it satisfies the hypothesis in Theorem 1.10. This equivalence has an analogue for manifolds of infinite volume stated by Gromov [GrHMGA, §4.1] which can be proved using similar techniques.
@@
Let $f:\mathbb{R}^{n}\to M$ be a wrapping map. Write $d\operatorname{vol}_{M}=\omega+d\alpha$, where $\omega$ is compactly supported and $\alpha$ is bounded. Then by Stokes’ theorem,
@@
Now let $K^{\prime}\supset K$ be another compact set which contains the $r$-neighborhood of $K$; we can assume $K^{\prime}$ is a smooth codimension zero submanifold of $M$ and give it a triangulation $\mathcal{T}$. Let $\lambda>0$ be such that every $\lambda$-ball in $K^{\prime}$ is contained in the star of some vertex of $\mathcal{T}$. Choose a triangulation $\mathcal{T}^{\prime}$ of $B(R)$ whose simplices are $C(n)$-bilipschitz to a linear simplex of side length $\min\{r,\lambda\}$ and have diameter $\min\{r/2,\lambda/2\}$. Then there is a subcomplex $X\subseteq\mathcal{T}^{\prime}$ which contains the image of $K$ and whose star lies in $K^{\prime}$. Moreover, by the usual proof of the simplicial approximation theorem, see e.g. [Hatc], a perturbation of $f$ gives a map $f^{\prime}$ whose restriction to $X$ is simplicial on $\mathcal{T}$.
We can assume that the induced map $\pi_{1}(K)\to\pi_{1}(M)$ is injective. Now consider the induced map from the dual graph of $X$ to the dual graph $\Gamma$ of $\widetilde{T}$, the universal cover of $T$. This defines a 0-cochain in $\Gamma$ of volume $\sim R^{n}$ whose coboundary has volume $\sim R^{n-1}$. Since $\Gamma$ is quasi-isometric to the Cayley graph of $\pi_{1}(K)$, it has the same isoperimetric and growth functions. Therefore, by the Varopoulos isoperimetric inequality (see e.g. [GrMS, §6.E]), $\pi_{1}(K)$ has growth rate $O(R^{n})$.
Now we could have taken $\pi_{1}(K)$ to be any finitely generated subgroup of $\pi_{1}(M)$. Therefore the nilpotency class of any such subgroup is at most n1n-1 (in fact, at most around $\sqrt{n}$, but this is harder to see), and so is the nilpotency class of $\pi_{1}(M)$. ∎
@@
First define a Lipschitz map of positive asymptotic degree $g_{n-2}:\mathbb{R}^{n-2}\to S^{n-2}$; for example, we can compose the universal covering map $\mathbb{R}^{n-2}\to T^{n-2}$ with a map that collapses the (n3)(n-3)-skeleton of the torus.
@@ Acknowledgements
The authors would like to thank Ben Lowe for comments drawing attention to the parallels between their previous work, which sparked this collaboration. F.M. would also like to thank Pekka Pankka and Susanna Heikkilä for useful conversations and comments, and Shmuel Weinberger for suggesting the proof of Theorem 1.6.